FAST ALGORITHMS FOR AB INITIO MOLECULAR DYNAMICS

\underline{JOHN M. HERBERT} and MARTIN HEAD-GORDON, Department of Chemistry, University of California, Berkeley, CA 94720.

We have developed a variety of novel algorithms for efficient propagation of classical molecular dynamics using electronic structure potentials, both at the self-consistent field level and beyond. One method, an extended Lagrangian algorithm, is inspired by Car-Parrninello molecular dynamics but employs atom-centered Gaussian orbitals rather than plane wave basis functions. This extends exisiting linear-scaling density functional theory algorithms, developed in our group, into the realm of ab initio molecular dynamics. An alternative set of techniques that accelerate molecular dynamics simulations employing post-Hartree-Fock electronic structure theory has also been developed, and can greatly expedite the fitting of accurate potential energy surfaces. We will survey these developments, as implemented in the Q-Chem electronic structure program.